OpenStudy (anonymous):

Does anyone have any special technique for remembering the entire unit circle? I have a memorization quiz on the entire think tomorrow and have most of all sine, cosine, tangent, and cotangent functions remembered out of quadrant 1. I still need help memorizing cosecant and secant. I know the trick that the denominator correlates back to the original ones in quadrant 1, but does anyone know any tricks to memorizing the numerators for each quadrant?

OpenStudy (zzr0ck3r):

there is a great way to do it with your hand, give me one sex

OpenStudy (zzr0ck3r):

err sec*

OpenStudy (zzr0ck3r):

http://www.youtube.com/watch?v=1-hrT1Ys390

OpenStudy (anonymous):

Thank you!

OpenStudy (zzr0ck3r):

np

OpenStudy (jhannybean):

LMAO @zzr0ck3r great typo.

OpenStudy (amistre64):

\[\frac{\sqrt{0,1,2,3,4}}{2}\]

OpenStudy (amistre64):

if you know sin and cos, the rest of them use these 2

OpenStudy (amistre64):

tan = sin/cos cot = cos/sin sec = 1/cos csc = 1/sin